40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Entry  Mon Jul 25 17:44:13 2016, Gabriele, Electronics, Characterization, High frequency noise budget qpd_noises.pngqpd_noise_budget.png
    Reply  Tue Jul 26 11:45:03 2016, Gabriele, Electronics, Characterization, High frequency noise budget noise_budget.pngnoise_budget.pngnoise_budget_whitened.png
Message ID: 61     Entry time: Mon Jul 25 17:44:13 2016     Reply to this: 62
Author: Gabriele 
Type: Electronics 
Category: Characterization 
Subject: High frequency noise budget 

I measured the noise sources limiting the QPD sensitivity. Unfortunately, I had to do some MATLAB tricks to get rid of the glitches: basically I load the data directly from the raw frames (NDS access to data is not working yet) and remove all jumps in the signals that happen in one single sample and are larger than a manually tuned threshold. This is not perfect, but it's enough to give us a rough idea of the spectrum of the QPD signal. The following plot shows the QPD_X signal (in units of disk motion, radians) in a few situations:

  • Blue: normal (laser on, room lights on)
  • Orange: laser on, room lights off (including the vacuum gauge)
  • Yellow: laser and room light off
  • Purple: same as above, but I switched off the PC monitors too
  • Green: QPD electronics off (this is the ADC noise)

The total power on the QPD is 30 uW, which correspond to a shot noise limited sensitivity of 4.3e-12 W/rHz. Considering that the signal is the quadrant asymmetry normalized by the total power, the shot noise limited sensitivity is sqrt(2) * SN / Power which once calibrated corresponds to 1.1e-10 rad/rHz.

The following plot shows that shot noise is the dominant source, followed closely by the electronics dark noise. The total agrees perfectly with the measured background noise above 2 kHz. Below that we have some leakage due to the large turbopump peak: this is due to FFT limitations but mostly to unsuppressed glitches.

From the QPD datasheet (Hamamatsu S5981) I learn that the noise equivalent power should be of the order of 2e-14 W/rHz at the sensitivity peak, so probably a factor of two or so worse at the HeNe frequency. It's still much lower than the measured dark noise. 

This sensitivity is already pretty good, but we can improve it by increasing the power on the diode. Indeed, 30 uW corresponds to about 2.7 V after the transimpedance, so we could increase the power by a factor 4 and win a factor 2 in the shot noise to dark noise ratio. Probably not worth it, since it will give us only a 30% gain in high frequency noise.

ELOG V3.1.3-